OpenStudy (anonymous):

Nine year old Chico is 1/5 of his aunt Sofie's age. How old is Aunt Sofie?

OpenStudy (anonymous):

\[c=1/5*aunt\]

OpenStudy (anonymous):

9 times by 5

OpenStudy (anonymous):

9=1 of 5 9*5=?

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let aunt's age is x, 9=1/5x x=9*5(when interchanging it will change its operations) =45 therefore aunt's age=45